OpenStudy (anonymous):

simplify the equation by combining like terms 4x+2x=12 i got six but cant figure out the rest

OpenStudy (anonymous):

6x = 12 x = 12/6 x = 2

OpenStudy (anonymous):

6x = 12 6x/6 = 12/6 x = 2